### v2

@@ Abstract
We address the problem of controlling the density of a large ensemble of follower agents by acting on a group of leader agents that interact with them. Using coupled partial integro-differential equations to describe leader and follower density dynamics, we establish feasibility conditions and develop two control architectures ensuring global stability. The first employs feed-forward control on the followers’ and a leaders’ density feedback. The second implements a dual feedback loop through a reference-governor that adapts the leaders’ density based on both populations measurements. Our methods, initially developed in a one-dimensional setting, are extended to multi-dimensional cases, and validated through numerical simulations for representative control applications, both for groups of infinite and finite size.
@@ 1 Introduction
Orchestrating the collective spatial organization of multi-agent systems is crucial in fields such as traffic control [1], collective additive manufacturing [2], synthetic biology [3], swarm robotics [4], and environmental management [5]. Leader-follower control strategies, where leader agents steer the behavior of a group of follower agents, are widely applied in these areas and other control applications [6, 7, 8, 9]. The use of controlled autonomous vehicles can improve traffic flows by avoiding stop-and-go waves and reducing emissions [10, 11], while in swarm robotics and synthetic biology, leader-follower dynamics facilitate large group management and cellular consortia regulation [12, 13].A critical challenge lies in establishing analytical guarantees for the achievement of desired collective tasks. For instance, this challenge is particularly evident in shepherding control problems, where a small group of controllable leader agents (or herders) must drive a larger population of follower agents toward a specified goal region. Various applications can be formulated as shepherding problems, including environmental pollutant management using robotic systems [5] and search and rescue operations [14]. Success in these applications critically depends on properly calibrating key parameters such as the ratio between leaders and followers, and their sensing capabilities [15, 16, 17].
@@ 1.1 Related work and main contributions
We also report significant work in deriving consensus-seeking strategies in networks of distributed parameter systems described by wave equations [41, 42][43].
@@ 2 Mathematical preliminaries
We denote with “ * ” the convolution operator. When referring to periodic domains and functions, the operator needs to be interpreted as a circular convolution [46]. We identify time and space partial differentiation with the subscripts $t$ and $x$, respectively. Integrals with explicit bounds are to be interpreted as indefinite (returning functions). Conversely, when the bounds are provided, they are definite integrals, returning constants, or, eventually, functions of the integral bounds. We denote by $\mathcal{C}^{p}(\mathcal{S})$, with $p\in\mathbb{N}_{\geq 0}$, the space of functions that are differentiable $p$ times with a continuous $p$-th derivative on $\mathcal{S}$. We use $W^{k,p}(\mathcal{S})$ for the Sobolev space of functions defined on $\mathcal{S}$ (the weak derivatives up to order $k$ are in $\mathcal{L}^{p}(\mathcal{S})$). Furthermore, when $p=2$, we use $H^{k}(\mathcal{S}):=W^{k,2}(\mathcal{S})$ [47].
@@
The inequality can be posed for more general bounded connected open domains $\Omega\subset\mathbb{R}^{d}$ with Lipschitz boundary [51, 47]. In this case, for any function $u\in W^{1,p}{\color[rgb]{0,0,1}(\Omega)}$ such that $\int_{\Omega}u\,\mathrm{d}\mathbf{x}=0$, the inequality reads
@@
When $p=2$, the optimal Poincaré constant for smooth bounded Lipschitz domains is $C=\tilde{d}/\pi$, where $\tilde{d}$ is least upper bound of the set of all distances between pairs of points in $\Omega$ [49] (for the domains we consider, $\tilde{d}=\pi$).
@@ 3 Model
$v^{FL}(x,t)=\int_{\mathcal{S}}f({\color[rgb]{0,0,1}y\triangleright x})\rho^{L}(y,t)\,\mathrm{d}y=(f*\rho^{L})(x,t)$ (19)
is a velocity field modeling the influence of the leaders on the dynamics of the followers, where ${\color[rgb]{0,0,1}y\triangleright x}=(x-y+\pi)\,\mathrm{mod}(2\pi)-\pi$ is the relative position between $x$ and $y$ wrapped on $\mathcal{S}$, and $f:\mathcal{S}\rightarrow\mathbb{R}$ is a soft-core (that is, returning bounded values when the argument is 0), odd interaction kernel [54]. To cope with the domain periodicity, we further assume $f$ to be periodic. Although the formulation is general, and any choice can be made for the kernel $f$, we fix it to be repulsive, that is, we choose
$f({\color[rgb]{0,0,1}x})=\frac{\mathrm{sgn}({\color[rgb]{0,0,1}x})}{\mathrm{exp}\left(\frac{2\pi}{L}\right)-1}\left[\mathrm{exp}\left(\frac{2\pi-|{\color[rgb]{0,0,1}x}|}{L}\right)-\mathrm{exp}\left(\frac{|{\color[rgb]{0,0,1}x}|}{L}\right)\right],$ (20)
@@
where $\rho^{F}_{0}(x)$ is periodic and such that $\int_{\mathcal{S}}\rho^{\color[rgb]{0,0,1}F}_{0}\,\mathrm{d}x=M^{F}$.
@@ 5.1 Example
$\hat{\rho}^{F}(x)=\frac{{\color[rgb]{0,0,1}\mathrm{exp}\left(\kappa\cos(x-\mu)\right)}}{2\pi I_{0}(\kappa)},$ (45)
@@ 6.1 Leaders’ control design
with $K_{L}>0$, makes the leaders’ error dynamics globally exponentially convergent to zero.
@@
$e^{L}(x,t)=e^{L}(x,0)\,\mathrm{exp}{\color[rgb]{0,0,1}(}-K_{L}t{\color[rgb]{0,0,1})}.$ (57)
@@ 7 Reference-governor control
To incorporate feedback related to the followers’ density, we introduce a reference-governor approach, inspired by [61]. We employ a dual-feedback loop structure: the outer loop, or governor loop, dynamically adjusts the target density for the leaders, $\hat{\rho}^{L}$, aiming to minimize the error $e^{F}=\bar{\rho}^{F}-\rho^{F}$ by facilitating the required organization of the followers. The inner loop, or the leaders’ control loop, then calculates the control input $u$ as per (18a), to guide the actual leaders’ density, $\rho^{L}$, towards $\hat{\rho}^{L}$, thereby reducing the discrepancy $e^{L}=\hat{\rho}^{L}-\rho^{L}$ to zero and fulfilling the control objectives. An illustration of this strategy is depicted in Fig. 3. Note that for controlling the leaders’ density we leverage the framework previously detailed in Section 6.1 (see Remark 9 specifically).
@@ 8 Numerical validation
We study a monomodal regulation problem. Specifically, we set $M^{L}=0.4$ and choose the von Mises distribution in (45) with $\kappa=1.8$ and $\mu=0$ for the desired followers’ density. We report the results of the numerical example in Fig. 4. Specifically, we show the initial and final displacement of the leaders’ and followers’ densities, resulting in the same steady-state profile with both the control techniques. Then, in Fig. 4(b) and 4(c) (upper panel), we report the time evolution of the percentage errors and KL divergences using respectively the feed-forward and the reference-governor control schemes. In Fig. 4(c) (bottom panel), we show the time evolution of the control function $\alpha$ selected according to (90). Similar results were obtained for bi-modal regulation tasks (see the Supplementary material available at [56].
@@ 9 An application to multi-agent leader-follower systems via continuification
$\mathrm{d}{x}_{k}^{F}$ $=\frac{1}{N^{L}+N^{F}}\sum_{j=1}^{N^{L}}f({\color[rgb]{0,0,1}x_{j}^{L}\triangleright x_{k}^{F}})\,{\color[rgb]{0,0,1}\mathrm{d}t}+\sqrt{2D}\mathrm{d}B_{k},$
@@
Contrary to the macroscopic simulations performed using the continuum formulation in Section 8, we register a small steady-state error in the discrete model. This error primarily arises from two factors: the finite size of the swarm, which, practically, constraints the continuum hypothesis to hold only partially, and the stochastic behavior of the followers.
@@ 10 Extension to higher dimensions
$\mathbf{v}^{FL}(\mathbf{x},t)=\int_{\Omega}\mathbf{f}({\color[rgb]{0,0,1}\mathbf{y}\triangleright\mathbf{x}})\rho^{L}(\mathbf{y},t)\,\mathrm{d}\mathbf{y}=(\mathbf{f}*\rho^{L})(\mathbf{x},t),$ (99)
@@ 10.1 Feasibility analysis
with periodic boundary conditions applied to $\mathbf{w}$ (we refer to [65] – Sec. 1.2.5, example 1.5 – for details about how the curl is computed in two-dimensions). Being $\mathbf{w}$ irrotational and $\Omega$ simply connected, we conclude that $\mathbf{w}=-\nabla\varphi$, where $\varphi$ is an unknown scalar potential.
@@
where $A$ is an arbitrary constant. This deconvolution of $\bar{\rho}^{L}$ is defined up to an arbitrary constant due to the linearity of the convolution operator and the assumption that the kernel is odd. Unlike the one-dimensional case, where $\mathbf{f}$ has a closed form, $H$ can only be computed numerically [60]. Consequently, the feasibility problem is reformulated in terms of the constant $A$.
@@ 10.2.2 Governor design
$\rho^{L}(\mathbf{x},t)=\hat{\rho}^{L}(x,t)+{\color[rgb]{0,0,1}\Lambda}(\mathbf{x},t),$ (110)
where $\hat{\rho}^{L}$ is the desired time-varying density and ${\color[rgb]{0,0,1}\Lambda}$ represents the transient behavior:
${\color[rgb]{0,0,1}\Lambda}(\mathbf{x},t)=-\left[\hat{\rho}^{L}(\mathbf{x},0)+\rho_{0}^{L}(\mathbf{x})\right]\mathrm{exp}(-K_{L}t).$ (111)
@@
$\mathbf{v}^{FL}(\mathbf{x},t)=\hat{\mathbf{v}}^{FL}(\mathbf{x},t)+(\mathbf{f}*{\color[rgb]{0,0,1}\Lambda})(\mathbf{x},t),$ (112)
@@
${\color[rgb]{0,0,1}\hat{\mathbf{v}}^{FL}}(\mathbf{x},t)=\bar{\mathbf{v}}^{FL}(\mathbf{x})+\alpha(t)\mathbf{w}(\mathbf{x},t),$ (113)
where $\bar{\mathbf{v}}^{FL}$ is derived from (103). Here, $\alpha{\color[rgb]{0,0,1}(t)}\in[0,1]$ is a control function to be determined, and
@@ Acknowledgments
$\hat{f}(x)=\mathrm{sgn}(x){\color[rgb]{0,0,1}\mathrm{exp}}\left(-\frac{\lvert x\rvert}{L}\right).$ (124)
@@
$f(x)=\sum_{k=-\infty}^{\infty}\mathrm{sgn}(x+2k\pi){\color[rgb]{0,0,1}\mathrm{exp}}\left({-\frac{\lvert x+2k\pi\rvert}{L}}\right).$ (125)
@@
$\phi(x)=(f*\rho)(x)=\frac{1}{\mathrm{e}^{\frac{2\pi}{L}}-1}\bigg{[}\mathrm{e}^{\frac{2\pi-x}{L}}\int_{-\pi}^{x}\mathrm{e}^{\frac{y}{L}}\rho(y)\,\mathrm{d}y\\ -\mathrm{e}^{\frac{x}{L}}\int_{-\pi}^{x}\mathrm{e}^{-\frac{y}{L}}\rho(y)\,\mathrm{d}y\\ {\color[rgb]{0,0,1}-}\mathrm{e}^{\frac{2\pi+x}{L}}{\color[rgb]{0,0,1}\int_{x}^{\pi}}\mathrm{e}^{-\frac{y}{L}}\rho(y)\,\mathrm{d}y{\color[rgb]{0,0,1}+}\mathrm{e}^{-\frac{x}{L}}{\color[rgb]{0,0,1}\int_{x}^{\pi}}\mathrm{e}^{\frac{y}{L}}\rho(y)\,\mathrm{d}y\bigg{]}.$ (126)
@@
$\rho(x)=\frac{1}{2}\int\left(\phi_{xx}-\frac{\phi}{L^{2}}\right)\,\mathrm{d}x+B,$ (128)

### v4

@@ Abstract
We address the problem of controlling the density of a large ensemble of follower agents by acting on a group of leader agents that interact with them. Using coupled partial integro-differential equations to describe leader and follower density dynamics, we establish feasibility conditions and develop two control architectures ensuring global stability. The first employs feed-forward control on the followers’ and a feedback on the leaders’ density. The second implements a dual feedback loop through a reference-governor that adapts the leaders’ density based on both populations’ measurements. Our methods, initially developed in a one-dimensional setting, are extended to multi-dimensional cases, and validated through numerical simulations for representative control applications, both for groups of infinite and finite size.
@@ 1 Introduction
Orchestrating the collective spatial organization of multi-agent systems is crucial in fields such as traffic control [1], collective additive manufacturing [2], synthetic biology [3], swarm robotics [4], and environmental management [5]. Leader-follower control strategies, where leader agents steer the behavior of a group of follower agents, are widely applied in these areas and other control applications [6, 7, 8, 9]. For instance, the use of autonomous vehicles can improve traffic flows by avoiding stop-and-go waves and reducing emissions [10, 11], while in swarm robotics and synthetic biology, leader-follower dynamics facilitate large group management and cellular consortia regulation [12, 13]. A critical challenge lies in establishing analytical guarantees for the achievement of desired collective tasks. This challenge is particularly evident in shepherding control problems, where a small group of controllable leader agents (or herders) must drive a larger population of follower agents toward a specified goal region. Various applications can be formulated as shepherding problems, including environmental pollutant management using robotic systems [5] and search and rescue operations [14]. Success in these applications critically depends on properly calibrating key parameters such as the ratio between leaders and followers, and their sensing capabilities [15, 16, 17].
@@ 1.1 Related work and main contributions
We also report significant work in deriving consensus-seeking strategies in networks of distributed parameter systems described by wave equations [41, 42, 43].
@@ 2 Mathematical preliminaries
We denote with “ * ” the convolution operator. When referring to periodic domains and functions, the operator needs to be interpreted as a circular convolution [46]. We identify time and space partial differentiation with the subscripts $t$ and $x$, respectively.Integrals with no explicit bounds are to be interpreted as indefinite (returning functions). Conversely, when the bounds are provided, they are definite integrals, returning constants, or, eventually, functions of the integral bounds. We denote by $\mathcal{C}^{p}(\mathcal{S})$, with $p\in\mathbb{N}_{\geq 0}$, the space of functions that are differentiable $p$ times with a continuous $p$-th derivative on $\mathcal{S}$. We use $W^{k,p}(\mathcal{S})$ for the Sobolev space of functions defined on $\mathcal{S}$ (the weak derivatives up to order $k$ are in $\mathcal{L}^{p}(\mathcal{S})$). Furthermore, when $p=2$, we use $H^{k}(\mathcal{S}):=W^{k,2}(\mathcal{S})$ [47].
@@
The inequality can be posed for more general bounded connected open domains $\Omega\subset\mathbb{R}^{d}$ with Lipschitz boundary [51, 47]. In this case, for any function $u\in W^{1,p}(\Omega)$ such that $\int_{\Omega}u\,\mathrm{d}\mathbf{x}=0$, the inequality reads
@@
When $p=2$, the optimal Poincaré constant for smooth bounded Lipschitz domains is $C=\tilde{d}/\pi$, where $\tilde{d}$ is least upper bound of the set of all distances between pairs of points in $\Omega$ [49] (for the periodic domains we consider, $\tilde{d}=\pi$).
@@ 3 Model
$v^{FL}(x,t)=\int_{\mathcal{S}}f({y\triangleright x})\rho^{L}(y,t)\,\mathrm{d}y=(f*\rho^{L})(x,t)$ (19)
is a velocity field modeling the influence of the leaders on the dynamics of the followers, where ${y\triangleright x}=(x-y+\pi)\,\mathrm{mod}(2\pi)-\pi$ is the relative position between $x$ and $y$ wrapped on $\mathcal{S}$, and $f:\mathcal{S}\rightarrow\mathbb{R}$ is a soft-core (that is, $|f(0)|<\infty$), odd interaction kernel [54]. To cope with the domain periodicity, we further assume $f$ to be periodic. Although the formulation is general, and any choice can be made for the kernel $f$, we fix it to be repulsive, that is,
$f({x})=\frac{\mathrm{sgn}({x})}{\mathrm{exp}\left(\frac{2\pi}{L}\right)-1}\left[\mathrm{exp}\left(\frac{2\pi-|{x}|}{L}\right)-\mathrm{exp}\left(\frac{|{x}|}{L}\right)\right],$ (20)
@@
where $\rho^{F}_{0}(x)$ is periodic and such that $\int_{\mathcal{S}}\rho^{F}_{0}\,\mathrm{d}x=M^{F}$.
@@ 5.1 Example
$\hat{\rho}^{F}(x)=\frac{{\mathrm{exp}\left(\kappa\cos(x-\mu)\right)}}{2\pi I_{0}(\kappa)},$ (45)
@@ 6.1 Leaders’ control design
with $K_{L}>0$, makes the leaders’ error dynamics globally point-wisely, exponentially convergent to zero.
@@
$e^{L}(x,t)=e^{L}(x,0)\,\mathrm{exp}\left(-K_{L}t\right).$ (57)
@@ 7 Reference-governor control
To incorporate feedback related to the followers’ density, we introduce a reference-governor approach, inspired by [61]. We employ a dual-feedback loop structure: the outer loop, or “governor loop”, dynamically adjusts the target density for the leaders, $\hat{\rho}^{L}$, aiming to minimize the error $e^{F}=\bar{\rho}^{F}-\rho^{F}$ by facilitating the required organization of the followers. The inner loop, or the leaders’ control loop, then calculates the control input $u$ as per (18a), to guide the actual leaders’ density, $\rho^{L}$, towards $\hat{\rho}^{L}$, thereby reducing the discrepancy $e^{L}=\hat{\rho}^{L}-\rho^{L}$ to zero and fulfilling the control objectives. An illustration of this strategy is depicted in Fig. 3. Note that for controlling the leaders’ density we leverage the framework previously detailed in Section 6.1 (see Remark 9 specifically).
@@ 8 Numerical validation
We study a monomodal regulation problem. Specifically, we set $M^{L}=0.4$ and choose the von Mises distribution in (45) with $\kappa=1.8$ and $\mu=0$ for the desired followers’ density. We report the results of the numerical example in Fig. 4. Specifically, we show the initial and final displacement of the leaders’ and followers’ densities, resulting in the same steady-state profile with both the control techniques. Then, in Fig. 4(b) and 4(c) (upper panel), we report the time evolution of the percentage errors and KL divergences using respectively the feed-forward and the reference-governor control schemes. In Fig. 4(c) (bottom panel), we show the time evolution of the control function $\alpha$ selected according to (90). Similar results were obtained for bi-modal regulation tasks (see the Supplementary material available at [56]).
@@ 9 An application to multi-agent leader-follower systems via continuification
$\mathrm{d}{x}_{k}^{F}$ $=\frac{1}{N^{L}+N^{F}}\sum_{j=1}^{N^{L}}f({x_{j}^{L}\triangleright x_{k}^{F}})\,{\mathrm{d}t}+\sqrt{2D}\mathrm{d}B_{k},$
@@
Contrary to the macroscopic simulations performed using the continuum formulation in Section 8, we register a small steady-state error in the discrete model. This error primarily arises from two factors: the finite size of the swarm, which, practically, constrains the continuum hypothesis to hold only partially, and the stochastic behavior of the followers.
@@ 10 Extension to higher dimensions
$\mathbf{v}^{FL}(\mathbf{x},t)=\int_{\Omega}\mathbf{f}(\mathbf{y}\triangleright\mathbf{x})\rho^{L}(\mathbf{y},t)\,\mathrm{d}\mathbf{y}=(\mathbf{f}*\rho^{L})(\mathbf{x},t),$ (99)
@@ 10.1 Feasibility analysis
with periodic boundary conditions applied to $\mathbf{w}$ (in two dimensions the curl operator returns a three dimensional vector with null third component – for more details see [65], Sec. 1.2.5, example 1.5). Being $\mathbf{w}$ irrotational and $\Omega$ simply connected, we conclude that $\mathbf{w}=-\nabla\varphi$, where $\varphi$ is an unknown scalar potential.
@@
where $A$ is an arbitrary constant. This deconvolution of $\bar{\rho}^{L}$ is defined up to an arbitrary constant due to the linearity of the convolution operator and the assumption that the kernel is odd and periodic. Unlike the one-dimensional case, where $\mathbf{f}$ has a closed form, $H$ can only be computed numerically [60]. Consequently, the feasibility problem is reformulated in terms of the constant $A$.
@@ 10.2.2 Governor design
$\rho^{L}(\mathbf{x},t)=\hat{\rho}^{L}(x,t)+{\Lambda}(\mathbf{x},t),$ (110)
where $\hat{\rho}^{L}$ is the desired time-varying density and ${\Lambda}$ represents the transient behavior:
${\Lambda}(\mathbf{x},t)=-\left[\hat{\rho}^{L}(\mathbf{x},0)+\rho_{0}^{L}(\mathbf{x})\right]\mathrm{exp}(-K_{L}t).$ (111)
@@
$\mathbf{v}^{FL}(\mathbf{x},t)=\hat{\mathbf{v}}^{FL}(\mathbf{x},t)+(\mathbf{f}*{\Lambda})(\mathbf{x},t),$ (112)
@@
${\hat{\mathbf{v}}^{FL}}(\mathbf{x},t)=\bar{\mathbf{v}}^{FL}(\mathbf{x})+\alpha(t)\mathbf{w}(\mathbf{x},t),$ (113)
where $\bar{\mathbf{v}}^{FL}$ is derived from (103). Here, $\alpha{(t)}\in[0,1]$ is a control function to be determined, and
@@ Acknowledgments
$\hat{f}(x)=\mathrm{sgn}(x){\mathrm{exp}}\left(-\frac{\lvert x\rvert}{L}\right).$ (124)
@@
$f(x)=\sum_{k=-\infty}^{\infty}\mathrm{sgn}(x+2k\pi){\mathrm{exp}}\left({-\frac{\lvert x+2k\pi\rvert}{L}}\right).$ (125)
@@
$\phi(x)=(f*\rho)(x)=\frac{1}{\mathrm{e}^{\frac{2\pi}{L}}-1}\bigg{[}\mathrm{e}^{\frac{2\pi-x}{L}}\int_{-\pi}^{x}\mathrm{e}^{\frac{y}{L}}\rho(y)\,\mathrm{d}y\\ -\mathrm{e}^{\frac{x}{L}}\int_{-\pi}^{x}\mathrm{e}^{-\frac{y}{L}}\rho(y)\,\mathrm{d}y\\ {-}\mathrm{e}^{\frac{2\pi+x}{L}}{\int_{x}^{\pi}}\mathrm{e}^{-\frac{y}{L}}\rho(y)\,\mathrm{d}y{+}\mathrm{e}^{-\frac{x}{L}}{\int_{x}^{\pi}}\mathrm{e}^{\frac{y}{L}}\rho(y)\,\mathrm{d}y\bigg{]}.$ (126)
@@
$\rho(x)=\frac{1}{2}\int\left(\phi_{xx}(x)-\frac{\phi(x)}{L^{2}}\right)\,\mathrm{d}x+B,$ (128)